40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5109     Entry time: Wed Aug 3 14:20:49 2011
Author: steve 
Type: Update 
Category: SUS 
Subject: ITMX and PRM damping restored 

ITMX watchdog tripped around 5 Torr  and the PRM around 450 Torr of this vent. They were restored than.

We are at 580 Torr now.

ELOG V3.1.3-